the present age of a mother is 30years older than her daughter.15years ago she was twice as old as her daughter.How old is the daughter at present?
Answer by ankor@dixie-net.com(22740)

the present age of a mother is 30 years older than her daughter.
m = d + 30
15 years ago she was twice as old as her daughter.
m - 15 = 2(d - 15)
replace m with (d+30)
(d+30) - 15 = 2(d - 15)
d + 15 = 2d - 30
15 + 30 = 2d - d
d = 45 is the daughter's age
:
:
Check this in the statement (m = 75)
"15 years ago she was twice as old as her daughter."
75 - 15 = 2(45-15)
60 = 2(30)